arie:

Sindhu successfully solved the system of equations 3x+y=9 and −2x−y=−7 . Which of these statements could correctly describe what she did? A She added the two equations together to solve for x, and then she substituted the value of x back into one of the equations and solved for y, getting x = 2 and y = 3. B She added the two equations together to solve for x, and then she substituted the value of x back into one of the equations and solved for y, getting x = 3 and y = 2. C She added the two equations together to solve for y, and then she substituted the value of y back into one of the equations and solved for x, getting x = 2 and y = 3. D She added the two equations together to solve for y, and then she substituted the value of y back into one of the equations and solved for x, getting x = 3 and y = 2.

lilyneverexisted:

A. She added the two equations together to solve for x, and then she substituted the value of x back into one of the equations and solved for y, getting =2x=2 and =3y=3.

jayfafr:

its d This option states that she added the two equations together to solve for y, which would eliminate x. This is correct. Then she substituted the value of y back into one of the equations and solved for x, which is the correct method. The solution states x = 3 and y = 2

wio:

It's A. By adding the two equations the result is: \[ (3x+y) + (-2x-y) = (9) + (-7) \\ 3x - 2x + y-y = 9 - 7 \\ x = 2 \] Substituting into one of the equations: \[ 3x + y = 9\\ 3(2) + y = 9\\ 6 + y = 9 \\ y = 9 - 6 \\ y = 3 \]